# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ VIII.1 Formation of post vacuum arc surface damage
Plasma built up filling gap between the electrodes, inevitably causes intensive heat exchange between the plasma and the metal surfaces of the electrodes. The heat exchange is mediated by the plasma particles, i.e. the ions and electrons as well as electromagnetic radiation, since the temperature of plasma can easily reach a few eV [6]. The mass difference between electrons and ions causes the reorganization of the plasma particles in such a way that the positively charged ions collect above the negatively charged surface forming a so-called Debye sheath [90] with the burning voltage of ( 20 eV) which drops over the distance of the order of a Debye length. An example of the potential drop near the surface, i.e. the sheath potential, can be seen in Figure 61 from Ref. [166]. The figure shows that this drop is not uniform and changes away from the center following the density distribution within the plasma itself.
@@
The ions escaping the plasma are accelerated via the electrostatic potential towards the surface causing a strong channel for the heat exchange along with ballistic collisions with surface atoms that also contribute to the surface morphology due to the plasma impact. In the early simulations of crater formation by plasma ions [164, 38], where the ideal circuit with the one-dimensional geometry of plasma was assumed to be the first-level approximation. In this approximation, the applied voltage between constituted the voltage drop in the Debye sheath without the loss. Although these simulations present rather academic interest, they already emphasize the ion flux effect on the crater shape. In Figure 62 one can see the signature features of the violent response of the surface material on the plasma impact. Although the experimental images are selected to show the side craters rather than the main cathode spot, even in these regions, the crater rims exhibit district shapes with well-visible fingers extended outside the rim. Simulations of thermally deposited energy of the same amount as the total energy deposited by the plasma ions resulted in crater shapes with smooth and round rims. This result manifests that the affected material at the cathode surface underwent a smooth transition from the solid to the molten phase in the absence of pressure imposed by plasma ions.
@@
Figure 63 shows a remarkable agreement between the crater profiles measured experimentally and in the simulations. Although the profiles were scaled 30-40 times to enable the comparison, the aspect ratio (the crater depth to crater diameter) is extremely well reproduced in simulations. For comparison, one crater profile obtained in the simulations of thermally deposited energy is added to emphasize the necessity of considering the directional momenta of plasma ions towards the cathode surface. This single profile is much shallower compared to those obtained in the MD simulations of plasma ion ”showers”.
@@
There were several attempts to simulate the formation of cathode spots in larger-scale simulations using the finite-element solvers of continuum equations. The macroscale of a typical cathode spot created by plasma on electrode surfaces motivates the description of plasma-surface interactions using continuous two-dimensional thermoplastic equations solved numerically [92, 110, 111]. In these simulations plasma is not seen as a shower of individual ions, but as a continuous pressure exerted on the surface within a radius roughly corresponding to the assumed lateral spread of plasma [162, 29, 28, 27]. These simulations can provide insights into the macroscopic behavior of metal surfaces simultaneously exposed to the pressure of plasma ion impacts and the heat exchange due to ion and electron currents, see an example shown in Figure 64, where the surface of anode was modified due to the plasma impact. The processes at the cathode are more violent and more accurate physical models where many ongoing processes can be included explicitly are needed. The present models are based on manually built-in assumptions, hence, they are limited in providing the first-principles insights into the plasma-surface interaction processes.
@@ IX.1 Introduction
All the physical processes of breakdown ignition described in the preceding sections - dislocation dynamics, field-biased surface diffusion, electron field emission and atomic thermal/field evaporation - are driven by the applied surface electric field and result in a nucleation site that is very small. The central role of the surface electric field and locality of breakdown initiation might lead one to expect that the limiting high-field performance of device would be determined solely by the maximum surface electric field in the device. And indeed over many decades, and for many devices, the maximum surface electric field has been used as a design constraint as given by the so-called Kilpatrick limit [78].
@@
In general terms, the observed dependence of peak surface electric field on device geometry is a consequence of the strength of the coupling of a electromagnetic energy of a system to the power absorbed by a developing breakdown arc. As a breakdown arc forms, it absorbs energy through the acceleration of emitted electrons resulting in a reduction of the fields around the breakdown site, slowing or stopping the evolution of the arc. Nearby stored energy and incoming power from the power source re-supplying the volume, either enough for the breakdown to fully form or if insufficient, then the arc extinguishes and is not observed in the macroscopic scale of the device. The coupling of the stored energy and supply power to a volume affected by an arc depends on the geometry of the system and the characteristics of the powering circuit. This coupling dependence means that the limiting performance of a system is given by a complex function of the electric and magnetic field properties of the system as a whole.
@@ IX.2 Phenomenological power-flow based models Implications for breakdown evolution
One of the earliest attempts to relate the field holding limit to the macroscopic characteristics of a high-field RF system, in this case an accelerating structure, is described in [1]. The focus of this work was to explain the difference in the level of surface damage caused by breakdowns between the more highly damaged high, and less damaged low group velocity ends of a tapered TW accelerating structures. Such structures were tested in the context of the NLC/JLC linear collider studies. The authors argued that the difference in damage was caused by a relative difference in the impedance match between a breakdown and the RF fields at that area of the structure. The difference in damage level was observed despite the near flat longitudinal profile of the peak surface electric-field on successive iris of these structures. The paper argues that a breakdown has a low impedance, it acts like a short circuit, This is then compared to the the impedance of the structure at the high group velocity end, where it is low, and the low group velocity end, where it is high. This results in a better impedance match, and consequently higher transfer of RF power, of the RF to an arc at the high-group velocity end compared to the low group velocity end. The group argued that the higher power coupling resulted in more damage to the structure surface. [1] quantifies this relationship by:
where is the power absorbed by the arc, is the local accelerating gradient, is the group velocity. relates the accelerating gradient to the stored energy per unit cell. Equation 5 is a proportionality for the absorbed power so does not give a specific limit that can be used in the design of an accelerating structure.
@@
where is the power flow, is the innermost circumference of the structure and is the empirically derived constant.
This limit is based on the general argument that the growth of an arc requires a certain level of power flow density in the volume occupied by the arc, and this density can be approximated by the quantity . This general argument was combined with trial comparisons of various functional dependencies to observed data from a series of about sixteen RF structures. The structures were made from roughly similar technology, from precision machined copper parts assembled by brazing or bonding and run in similar conditions. The quantity was most consistent with the general argument and with the available data. By basing the value of the limit on a set of tested structures, , can then be used to predict the achievable gradient in a new design, but using similar technology and operated similarly.
The power flow density is approximated by the input power divided to the circumference of the inner surface rather than the cross section because the power flow in the disk loaded waveguide geometry typically used in accelerating structures is confined to the near surface of the inner radius of the iris. The height of the power flow region is similar to that of a typical plasma spot thus the the quantity describes this power density reasonably well.
@@
The limit is consistent with the observation that an arc’s impedance evolves to match the power source impedance. The adaptability of arc impedance to drive impedance as is see in DC arcs where the impedance of an arc adjusts itself so that the burning voltage goes to the ionization energy of the electrode material, around 20 V in the case of copper [5]. This focus on power density rather than impedance matching is one major difference compared to using 5 as a high-field limit.
@@
One is that that predicts very well the high-field limit for TW structures of a given frequency. This is true over a wide range of structure types from accelerating structures with a group velocity of around 10 % to a power generating structures with a group velocity of around 50 %. It is likely that 5 would fit the same constant frequency TW subsets of data equally well since a larger aperture structure has a higher group velocity for the same phase advance per cell.
@@
On the other hand, clearly does not apply to SW structures. There is no net power flow in such structures which implies an infinite field holding capability. 5 fails in the same way.
@@
In addition, the frequency dependence of implies that achievable gradient for a scaled geometry should go up as frequency goes down, which is in contradiction to the the observation that gradient is approximately constant for scaled structures. Indeed, some early publications have concluded that gradient goes up with frequency. This is addressed in [182] and resolved by explicitly adding a proportionality to frequency dependence. This is equivalent to the statement that achievable gradient is independent of frequency for a scaled geometry. It is also the flat frequency dependence predicted by 5.
@@
Finally neither 6 nor 5 capture the pulse length dependence observed in short, sub-s pulses. We have seen in section III.5 that a pulse length dependence, lower field holding for longer pulses, is only observed in pulse lengths up to around 1 s after which field holding remains constant. The models that will be derived below also do not capture the pulse length dependence. It is likely that this is because the mechanism that gives the pulse length dependence does not come from circuit and power flow related effects. The RF and DC voltage pulses described in this report have minimum duration of the order of 100 ns. On the other hand the turn-on time for breakdowns is much shorter, of the order of ns as seen in the pulse shape data shown in section III and the simulation results shown in section VII. The pulse length dependence is more likely to come from dislocation processes as described in section IV. The origin of pulse length dependence is more likely to be from dislocation dynamics as described in IV.
@@
The next generation high-field limit, referred to as , extends the concept that fields are limited by the level of the local power flow at the position of igniting arc. The quantity was introduced in the publication [58]. The publication begins a greatly expanded compilation of 30 Ghz and X-band testing results compared to [182]. The publication [58] includes subsequent structures tested in the context of the linear collider studies compared to [182], but also includes more details of the tests including measured breakdown rates as well as field quantities such as peak , peak and others re-simulated from structure geometries collected from the groups who built the structures and carried out the tests. Although many differences between different structure tests remain, this broader and deeper compilation of experimental results provides an improved basis for evaluating the applicability of high-gradient limits.
@@
The quantity is based on the Poynting vector, so is explicitly related to the local power flow inside an RF structure. It is similar to in the sense that it is related to power flow, but different in that it is calculated locally. is a scalar quantity that combines the real part of the Poynting vector with fraction of the imaginary part of the Poynting vector. The real part of the Poytning vector is the net power flow, while the imaginary part is the power flow that occurs within an RF cycle as energy is transferred back and forth between high electric and magnetic field regions of an RF structure. captures the fact that both flows of energy contribute to maintaining fields in the local region of an arc. The quantity is:
@@
The power flowing at a specific point on the surface is available to provide power to the processes driving the evolution of an arc. In the publication, the power is found to be comparable to the power necessary to heat a field emitting tip. The value of the fraction is determined by a phenomenological fit and is explained by relative coupling of the real and imaginary power flow to the tip. A schematic image of power flow near a field emitting tip is shown in Fig 65.
@@
The inclusion of imaginary power flow is crucial for the applicability of to SW structures. SW structures do not have a real power flow (strictly speaking they have a power flow related to replenishing power loss due to finite quality factory, but this is extremely low). As we have seen this means that the and limits would imply infinite field levels. SW structures do however have imaginary power flow, the power flow that occurs withing a each cycle of RF fields, which is available to power a breakdown. This power is captured in the second term in 7. The applicability of to SW structures resolves a major deficiency of the previous limits.
@@
It has also been suggested that gradient is limited by surface magnetic field, based on results of a particular set of structures with different irises [40]. However, there does not appear to be a numerical limit generally applicable across a broad range of devices. Also in the particular structures of [40], other quantities such as are often also constant. Also looking at tests of structures beyond the group, for example tests of HOM damped and undamped CLIC structures show that there is of the order of difference in achievable accelerating gradient and similar values of peak surface electric field and while there is a difference of more than a factor of two in surface magnetic field. Magnetic field is indeed relevant for the pulsed surface heating limit, but this is not a breakdown limit.
@@ IX.3 Power coupling between the macroscopic system and the microscopic arc development
We have seen that successive generations of high-field limits have been capable of describing an increasing range of RF structures; gone from describing performance based on global characteristics only to capturing the effect of local features and providing an increasingly detailed physical picture of the breakdown initiation process.
@@
None of the three RF limits, -based, , and , however, are able to describe a DC system, like that described in II.3 or other high-voltage DC systems, because there is no power flow - . There is a power flow during charging and discharging of the electrode capacitance in the DC system, but the value is extremely low compared to typical values in RF structures. In addition, breakdowns are distributed in time throughout the high voltage pulse, not only during the rise and fall. On the other hand, a clear dependency of the surface electric field limit is observed for the pulsed DC system in analogy to the geometrical dependency observed in RF systems and described by -based, , and . The dependency observed in the DC system is that the maximum voltage is proportional to the gap between the electrodes to the power of 0.7, with the consequence that the peak surface electric field goes down as the electrode gap increases. This 0.7 dependence can be seen in varying gap data taken with the pulsed DC system as shown in Fig. 26.
@@
A specific discrepancy with an experiment of emerged from an RF structure which was tested after the publication of [58]. The structure was a so-called crab cavity, also know as a deflecting structure. The fundamental difference to previously tested high-gradient structures is that the crab cavity has an angularly varying field pattern, it operates in the TM110 mode in contrast to the circularly symmetric, non-angularly varying pattern of the previously tested and discussed structures which all operate in the TM010 mode. The field pattern is shown in Fig. 67. The structure was high gradient tested and cut open after test. The breakdown locations from the entirety of the test were determined by optical imaging. Breakdowns are seen as dots in Fig. 67. One can clearly see that the breakdown density is highest where the surface electric field is highest. The breakdowns are not concentrated where is highest, not surface electric field. Surprisingly the structure operated at the standard value of X W/, but the lack of correlation between regions of high and high breakdown density indicates that is not a sufficient quantity as a general high gradient limit. Neither -based not limits make any statement about where inside a structure breakdowns are likely to occur.
@@
In addition to the discrepancy of with the experimental result of the crab cavity, a further theoretical concern became increasingly clear that also applies to the -based and high-field limits. This is that all three quantities are functions of single-frequency, steady-state solutions of the electromagnetic fields. On the other hand, breakdown is a very fast phenomenon, occurring at a ns timescale as we have seen. Since the high-field limits are meant to capture the threshold of breakdown, a more comprehensive limit must capture the fast time response of the electromagnetic fields induced by the rapidly evolving breakdown. The focus of this report is on the initiation of breakdown, so the earliest moments of the breakdown process become critical. The consequence is that a more broadband analysis of the high-field devices is necessary.
@@
The generalization from earlier high-field limits is how much power is available to flow into the small volume containing the breakdown in the short, ns-scale, time while the breakdown is evolving from quiescence, and whether this is sufficient or insufficient to drive the breakdown evolution process. This generalization is crucial to covering DC high-field systems since it moves from steady-state power flow, 0 in the case of a DC system, to the transient flow of power into the breakdown area. In the case of both DC and RF, the power is available from locally stored energy on short timescales, the speed of light in free space is 30 cm/ns to give an approximate scale, and from the power circuit on longer timescales.
@@
Another way of describing this concern is posing the question will the growing electron emission from the breakdown site load the local fields enough that the field drops below the values needed to sustain the physical processes described in section VII. To this end, a next-generation high-field limit has been developed. It seeks to consider all data and results as previous limits and, in addition, incorporate the newest data and considerations just described. Furthermore, it is based on a rigorous formulation of the power coupling dynamics that incorporates the transient electromagnetic phenomena that occur during breakdown. The fundamental aspects of this formulation were introduced by J. Paszkiewicz [136], but here we shall present it in a slightly different and more general manner.
@@
Due to the linearity of the Maxwell equations, can be considered as a superposition of the unloaded field originating from the external power sources and a loading field caused by the moving charges of the arc itself. In principle, depends on the entire distribution of the emitted current density in a system, i.e., the entire trajectories of electrons and ions emitted by the arc. However, we can approximate that at the breakdown site depends solely on the current density distribution in its close vicinity, as the contribution of farther moving charges decays rapidly due to two effects: the moving charge dispersion that reduces the current density and the dropping contribution of charges with increasing distance, in view of the generalized Coulomb’s law (Jefimeko’s electromagnetic field equations [57]).
@@
has units of , whose physical meaning becomes apparent by multiplying eq. (9) by , and defining , and . Thus we get the dependence that motivates the circuit structure shown in Figure 57. This analysis reduces the entire electromagnetic coupling of the macroscopic system to the arc into a simple local impedance function. The term scales the impedance with respect to the length of the arc domain, where the presence of moving charges is considered.

### v2

@@ VIII.1 Formation of post vacuum arc surface damage
Plasma built up that fills the gap between the electrodes inevitably causes intensive heat exchange between the plasma and the metal surfaces of the electrodes. Heat exchange is mediated by plasma particles, that is, ions, electrons and electromagnetic radiation, since the temperature of the plasma can easily reach a few eV [6]. The mass difference between electrons and ions causes reorganization of plasma particles in such a way that positively charged ions accumulate above the negatively charged surface, forming a so-called sheath [95] with a burning voltage of 20 eV which drops over a distance of the order of a Debye length. An example of a potential drop near the surface, that is, the sheath potential, can be seen in Fig. 62. The figure shows that this drop is not uniform and changes away from the center following the density distribution within the plasma itself.
@@
The ions escaping the plasma are accelerated via the electrostatic potential towards the surface, causing a strong channel for the heat exchange along with ballistic collisions with surface atoms that also contribute to the surface morphology as a result of the plasma impact. In early simulations of crater formation by plasma ions [174, 41], the ideal circuit with the one-dimensional geometry of the plasma was assumed as the first-level approximation. In this approximation, the applied voltage to the electrodes constituted the voltage drop in the sheath without loss. Although these simulations are of rather academic interest, they already emphasize the ion flux effect on the crater shape. From the visual agreement between the simulation and the experimental images shown in Fig. 63, one can see the signature features of the violent response of the surface material to the plasma impact. Although the experimental images are selected to show the side craters rather than the main cathode spot, one clearly sees that even in these regions, the crater rims exhibit distinct shapes with well-visible fingers extended outside the rim.
In the simulations where the same amount of energy as deposited in the plasma impact simulation was instantaneously deposited as thermal energy ( 10 eV/atom) within a cylinder of the same radius as the plasma impact spot, the formed crater was fully symmetric with smooth rounded rims and no apparent asperities. The result indicates a very fast but still uniform phase transition from solid to molten, and to vapor. No overlapping impacts of plasma ions took place in these simulations, hence, no thermal instabilities responsible for asymmetric shape of rims in the simulations shown in Fig. 63.
@@
Fig. 64 shows a remarkable agreement between the crater profiles measured experimentally and in the simulations of plasma impacts. Although the profiles were scaled 30-40 times to enable comparison, the aspect ratio (crater depth to crater diameter) is extremely well reproduced in simulations. For comparison, one crater profile obtained in the simulations of thermally deposited energy is added to emphasize the necessity of considering the directional momenta of plasma ions towards the cathode surface. This single profile is much shallower than those obtained in the MD simulations of plasma ion showers, which results from the lack of plasma pressure on the cathode spot due to energetic plasma ions accelerated toward the surface.
@@
There were several attempts to simulate the formation of cathode spots in larger-scale simulations using the finite-element solvers of continuum equations. The macroscale of a typical cathode spot created by plasma on electrode surfaces motivates the description of plasma-surface interactions using continuous two-dimensional thermoplastic equations solved numerically [97, 118, 120]. In these simulations, plasma is not seen as a shower of individual ions, but as a continuous pressure exerted on the surface within a radius roughly corresponding to the assumed lateral spread of plasma [172, 31, 30, 29]. These simulations can provide insights into the macroscopic behavior of metal surfaces simultaneously exposed to the pressure of plasma ion impacts and the heat exchange due to ion and electron currents, an example is shown in Fig. 65, where the surface of the anode was modified due to the plasma impact. The processes at the cathode are more violent and more accurate physical models where many ongoing processes can be included explicitly are needed. The present models are based on manually built-in assumptions, hence, they are limited in providing the first-principles insights into the plasma-surface interaction processes.
@@ IX.1 Introduction
All the physical processes of breakdown ignition described in the preceding sections - dislocation dynamics, field-biased surface diffusion, electron field emission and atomic thermal/field evaporation - are driven by the applied surface electric field and result in a very small nucleation site. The central role of the surface electric field and locality of breakdown initiation might lead one to expect that the limiting high-field performance of a device would be determined solely by the maximum surface electric field in the device. And indeed over many decades, and for many devices, the maximum surface electric field has been used as a design constraint as given by the so-called Kilpatrick limit [82].
@@
In general terms, the observed dependence of peak surface electric field on device geometry is a consequence of the strength of the coupling of electromagnetic energy of a system to the power absorbed by a developing breakdown arc. As a breakdown arc forms, it absorbs energy through the acceleration of emitted electrons, resulting in a reduction of the fields around the breakdown site, slowing or stopping the evolution of the arc. Nearby stored energy and incoming power from the power source re-supply the volume, either enough for the breakdown to fully form or, if insufficient, then the arc extinguishes and is not observed in the macroscopic scale of the device. The coupling of the stored energy and supply power to a volume affected by an arc depends on the geometry of the system and the characteristics of the powering circuit. This coupling dependence means that the limiting performance of a system is given by a complex function of the electric and magnetic field properties of the system as a whole.
@@ IX.2 Phenomenological power-flow based models Implications for breakdown evolution
One of the earliest attempts to relate the field holding limit to the macroscopic characteristics of a high-field rf system, in this case an accelerating structure, is described in [1]. The focus of this work was to explain the difference in the level of surface damage caused by breakdowns between the more highly damaged higher, and less damaged lower-group velocity ends of a tapered traveling wave accelerating structures. Such structures were tested in the context of the NLC/JLC linear collider studies. The authors argued that the difference in damage was caused by a relative difference in the impedance match between a breakdown and the rf fields at that area of the structure. The difference in damage level was observed despite the near flat longitudinal profile of the peak surface electric-field on successive iris of these structures. The paper argues that a breakdown has a low impedance and it acts like a short circuit. This is then compared to the the impedance of the structure at the high group velocity end, where it is low, and the low group velocity end, where it is high. This results in a better impedance match, and consequently higher transfer of rf power to an arc at the high-group velocity end compared to the low group velocity end. The group argued that the higher power coupling resulted in more damage to the structure surface. [1] quantifies this relationship by:
where is the power absorbed by the arc, is the local accelerating gradient, is the group velocity and relates the accelerating gradient to the stored energy per unit cell. Equation (7) is a proportionality for the absorbed power so does not give a specific limit that can be used in the design of an accelerating structure.
@@
where is the power flow, is the innermost circumference of the structure and is an empirically derived constant.
This limit is based on the general argument that the growth of an arc requires a certain level of power flow density in the volume occupied by the arc, and this density can be approximated by the quantity . This general argument was combined with trial comparisons of various functional dependencies to observed data from a series of about sixteen rf structures. The structures were made from roughly similar technology, from precision machined copper parts assembled by brazing or bonding and run in similar conditions. The quantity was most consistent with the general argument and with the available data. By basing the value of the limit on a set of tested structures, can then be used to predict the achievable gradient in a new design, but using similar technology and operated similarly.
The power flow density is approximated by the input power divided to the circumference of the inner surface rather than the cross section because power flow in the disk loaded waveguide geometries typically used in accelerating structures is confined to the near the surface of the inner radius of the iris. The height of the power flow region is similar to that of a typical plasma spot thus the the quantity describes this power density reasonably well.
@@
The limit is consistent with the observation that an arc’s impedance evolves to match the power source impedance. The adaptability of arc impedance to drive impedance is seen in dc arcs where the impedance of an arc adjusts itself so that the burning voltage goes to a material-specific value that depends on its ionization levels, around 20 V in the case of copper [5]. This focus on power density rather than impedance matching is one major difference compared to using Eq. (7) as a high-field limit.
@@
One is that predicts very well the high-field limit for traveling wave structures of a given frequency. This is true over a wide range of structure types, from accelerating structures with a group velocity of around 10 % to power-generating structures with a group velocity of around 50 %. It is likely that Eq. (7) would fit the same constant frequency traveling wave subsets of data equally well since a larger aperture structure has a higher group velocity for the same phase advance per cell.
@@
On the other hand, clearly does not apply to standing wave structures. There is no net power flow in such structures which implies an infinite field holding capability. Equation (7) fails in the same way.
@@
In addition, the frequency dependence of implies that the achievable gradient for a scaled geometry should go up as frequency goes down, which is in contradiction to the observation that the gradient is approximately constant for scaled structures. Indeed, some early publications have concluded that the gradient goes up with frequency. This is addressed in [192] and resolved by explicitly adding a proportionality to frequency dependence. This is equivalent to the statement that an achievable gradient is independent of frequency for a scaled geometry. It is also the flat frequency dependence predicted by Eq. (7).
@@
Finally, neither Eq. (8) nor Eq. (7) capture the pulse length dependence observed in short, sub-s pulses. We have seen in section III.5 that a pulse length dependence is only observed in pulse lengths up to around 1 s after which field holding remains constant. The models that will be derived below also do not capture the pulse length dependence. This is likely because the mechanism that gives the pulse length dependence does not come from circuit and power flow-related effects. The rf and dc voltage pulses described in this report have a minimum duration of the order of 100 ns. On the other hand the turn-on time for breakdowns is much shorter, of the order of ns as seen in the pulse shape data shown in section III and the simulation results shown in section VII. The pulse length dependence is more likely to come from dislocation processes as described in section IV.
@@
The next generation high-field limit, referred to as , extends the concept that fields are limited by the level of the local power flow at the position of the igniting arc. The quantity was introduced in [61]. The publication begins a greatly expanded compilation of 30 GHz and X-band testing results compared to [192]. and also includes more details of the tests including measured breakdown rates as well as field quantities such as peak , peak and others re-simulated from structure geometries collected from the groups who built the structures and carried out the tests. Although many differences between different structure tests remain, this broader and deeper compilation of experimental results provides an improved basis for evaluating the applicability of high-gradient limits.
@@
The quantity is based on the Poynting vector, so is explicitly related to the local power flow inside an rf structure. It is similar to in the sense that it is related to power flow, but different in that it is calculated locally. is a scalar quantity that combines the real part of the Poynting vector with fraction of its imaginary part. The real part of the Poytning vector is the net power flow, while the imaginary part is the power flow that occurs within an rf cycle as energy is transferred back and forth between high electric and magnetic field regions of an rf structure. captures the fact that both flows of energy contribute to maintaining fields in the local region of an arc. The quantity is:
@@
The power flowing at a specific point on the surface can power the processes driving the evolution of an arc and is found to be comparable to the power necessary to heat a field-emitting tip. The value of the fraction is determined by a phenomenological fit and is explained by the relative coupling of the real and imaginary power flow to the tip. A schematic image of power flow near a field-emitting tip is shown in Fig. 66.
@@
The inclusion of imaginary power flow is crucial for the applicability of to standing wave structures. Standing wave structures do not have a real power flow (strictly speaking they do have a power flow related to replenishing power loss due to finite quality factors, but this is extremely low). As we have seen this means that the and limits would imply infinite field levels. Standing wave structures do however have imaginary power flow, the power flow that occurs within each cycle of rf fields, which is available to power a breakdown. This power is captured in the second term in section Eq. (9). The applicability of to standing wave structures resolves a major deficiency of the previous limits.
@@
It has also been suggested that gradient is limited by surface magnetic field, based on results of a particular set of structures with different irises [43]. However, there does not appear to be a numerical limit generally applicable across a broad range of devices. Also in the particular structures of [43], other quantities such as are often also constant. Also looking at tests of structures beyond the group, for example tests of higher-order-mode damped and undamped CLIC structures show that there is of the order of difference in achievable accelerating gradient and similar values of peak surface electric field and while there is a difference of more than a factor of two in surface magnetic field. Magnetic field is indeed relevant for the pulsed surface heating limit, but it is not a breakdown limit.
@@ IX.3 Power coupling between the macroscopic system and the microscopic arc development
We have seen that successive generations of high-field limits have been capable of describing an increasing range of rf structures; from describing performance based on global characteristics only, to capturing the effect of local features and providing an increasingly detailed physical picture of the breakdown initiation process.
@@
None of the three rf limits, -based, , and , however, are able to describe a dc system, like that described in Sec. II.3 or other high-voltage dc systems, because there is no power flow (). There is a power flow during charging and discharging of the electrode capacitance in the dc system, but the value is extremely low compared to typical values in rf structures. In addition, breakdowns are distributed in time throughout the high voltage pulse, not only during the rise and fall. On the other hand, a clear dependency of the surface electric field limit is observed for the pulsed dc system in analogy to the geometrical dependency observed in rf systems and described by -based, , and . The dependency observed in the dc system is that the maximum voltage is proportional to the gap between the electrodes to the power of 0.7, with the consequence that the peak surface electric field goes down as the electrode gap increases. This 0.7 dependence can be seen in varying gap data taken with the pulsed dc system as shown in Fig. 28.
@@
A specific discrepancy with an experiment of emerged from an rf structure which was tested after the publication of [61]. The structure was a so-called crab cavity, also know as a deflecting structure. The fundamental difference to previously tested high-gradient structures is that the crab cavity has an angularly varying field pattern. It operates in the TM110 mode in contrast to the circularly symmetric, non-angularly varying pattern of the previously tested and discussed structures which all operate in the TM010 mode. The field pattern is shown in Fig. 68. The structure was high gradient tested and cut open after test. The breakdown locations from the entirety of the test were determined by optical imaging. Breakdowns are seen as dots in Fig. 68. One can clearly see that the breakdown density is highest where the surface electric field is highest, not concentrated at the location of maximum nor the location of maximum surface magnetic field. Surprisingly, the structure operated at the typical value of , but the lack of correlation between regions of high and high breakdown density indicates that is not a sufficient quantity as a general high gradient limit. Neither -based nor limits make any statement about where inside a structure breakdowns are likely to occur.
@@
In addition to the discrepancy of with the experimental result of the crab cavity, a further theoretical concern became increasingly clear which also applied to the -based and high-field limits. This is that all three quantities are functions of single-frequency, steady-state solutions of the electromagnetic fields whereas breakdown is a very fast, ns timescale, phenomenon. Since the high-field limits are meant to capture the threshold of breakdown, a more comprehensive limit must capture the fast time response of the electromagnetic fields induced by the rapidly evolving breakdown. The focus of this review is on the initiation of breakdown, so the earliest moments of the breakdown process become critical. The consequence is that a more broadband analysis of the high-field devices is necessary.
@@
The generalization from earlier high-field limits is how much power is available to flow into the small volume containing the breakdown in the short, ns time-scale when the breakdown is evolving from quiescence, and whether this is sufficient or insufficient to drive the breakdown evolution process. This generalization is crucial to covering dc high-field systems since it moves from steady-state power flow to the transient flow of power into the breakdown area. In the case of both dc and rf, the power is available from locally stored energy on short timescales (the speed of light in free space is 30 cm/ns to give an approximate scale) and from the power circuit on longer timescales.
@@
Another way of describing this concern is posing the question will the growing electron emission from the breakdown site load the local fields enough that the field drops below the values needed to sustain the physical processes described in section VII. To this end, a next-generation high-field limit has been developed. It seeks to consider all data and results as previous limits and, in addition, incorporate the newest data and considerations just described. Furthermore, it is based on a rigorous formulation of the power coupling dynamics that incorporates the transient electromagnetic phenomena that occur during breakdown. The fundamental aspects of this formulation were introduced by [145], but here we shall present it in a slightly different and more general manner.
@@
Due to the linearity of the Maxwell equations, can be considered as a superposition of the unloaded field originating from the external power sources and a loading field caused by the moving charges of the arc itself. In principle, depends on the entire distribution of the emitted current density in a system, i.e. the entire trajectories of electrons and ions emitted by the arc. However, we can approximate that at the breakdown site depends solely on the current density distribution in its close vicinity, as the contribution of farther moving charges decays rapidly due to two effects: the moving charge dispersion that reduces the current density and the dropping contribution of charges with increasing distance, in view of the generalized Coulomb’s law (Jefimeko’s electromagnetic field equations [60]).
@@
has units of , whose physical meaning becomes apparent by multiplying Eq. (11) by , and defining , and . Thus we get the dependence that motivates the circuit structure shown in Fig. 58. This analysis reduces the entire electromagnetic coupling of the macroscopic system to the arc into a simple local impedance function. The term scales the impedance with respect to the length of the arc domain, where the presence of moving charges is considered.